H8 Q7 Suppose that, prior to the passage of the Truth in Lending Simplification Act and...

H8 Q7

Suppose that, prior to the passage of the Truth in Lending Simplification Act and Regulation Z, the demand for consumer loans was given by Qdpre-TILSA = 10 -80P (in billions of dollars) and the supply of consumer loans by credit unions and other lending institutions was QSpre-TILSA = 4 + 120P (in billions of dollars). The TILSA now requires lenders to provide consumers with complete information about the rights and responsibilities of entering into a lending relationship with the institution, and as a result, the demand for loans increased to Qdpost-TILSA = 16 -80P (in billions of dollars). However, the TILSA also imposed “compliance costs” on lending institutions, and this reduced the supply of consumer loans to QSpost-TILSA = 2 + 120P (in billions of dollars).

Based on this information, compare the equilibrium price and quantity of consumer loans before and after the Truth in Lending Simplification Act.(Note: Q is measured in billions of dollars and P is the interest rate).

Instruction: Enter your responses for the equilibrium price in percentage terms, and round all responses to one decimal place.

Equilibrium price (interest rate) before TILSA: percent

Equilibrium quantity (in billions of dollars) before TILSA: $  billion

Equilibrium price (interest rate) after TILSA:  percent

Equilibrium quantity (in billions of dollars) after TILSA: $ billion
Answer #1

At equilibrium we have

10 – 80P = 4 + 120P

6 = 200P

P = 0.03 or 3%

Q = 10 – 2.4 = 7.6 billion of dollars

New demand Qdpost-TILSA = 16 -80P (in billions of dollars) and new supply of consumer loans QSpost-TILSA = 2 + 120P (in billions of dollars).

16 – 80P = 2 + 120P

14 = 200P

P = 7% or 0.07

Q = 10.4 billion of dollars

Hence,
Equilibrium price (interest rate) before TILSA: 3 percent

Equilibrium quantity (in billions of dollars) before TILSA: $7.6 billion

Equilibrium price (interest rate) after TILSA:  7 percent

Equilibrium quantity (in billions of dollars) after TILSA: $10.4 billion